OpenStudy (anonymous):

if r=3 and x=7 then what is the volume of the cylinder

OpenStudy (anonymous):

The volume is the area of one of the circles (top/bottom) times the height. Area of a circle = pi*r^2 V = h(pi*r^2) V = h(pi*3^2) V = h(9pi) You didn't specify a height or where x fits into anything, so that's all I have.